DO A QUADRATIC FIT ON THE POSITION VS, TIME GRAPH, THE VALUE OF THE QUADRATIC A QUAL TO HALF OF THE ACCELERATION. WHY? D A ALUINEAR FIT ON THE VELOCITY VS. TIME GRAPH. THE VALUE OF THE SLOPE IS EQUAL TO THE ACCELERATION, WHY? COMPARE THE ACCELERATION VALUES FROM BOTH GRAPHS. DO PERCENT DIFFERENCE. POST-LAB QUESTIONS: IF THE CART DID NOT ACCELERATE, WHAT WOULD BE THE SHAPES OF THE POSITION VS. TIME AND THE VELOCITY VS. TIME GRAPHS? WHY? MAKE SKETCHES OF THE 2 GRAPHS. 0.4 Eman sies, Sanart a.

need help please
Answer #1

If a particle starts from position x_0 with initial velocity u , and acceleration a (constant).

Position of particle as a function of time is x=x_0+ut+rac{1}{2}at^2

During the position time plot(quadratic fit), equation used is x=At^2+Bt+C

Comparing the two equations, C=x_0,B=u,A=rac{1}{2}a

Hence A is equal to half of the acceleration.

---------------

When a particle moves with initial velocity u , and acceleration a

Velocity of a particle as a function of time is v=u+at

For the linear fit of velocity Vs time, the equation used is v=mt+b

comparing the above two equations, m=a, b=u

That is slope of best fit line m is equal to acceleration (a ).

-----------------

Acceleration from quadratic fit is A=rac{1}{2}a_1=0.0731Rightarrow a_1=0.1462,m/s^2

Acceleration from best fit line is a_2=m=0.152,m/s^2

Percentage difference in accelerations is rac{|a_1-a_2|}{a_2}*100=rac{|0.1462-0.152|}{0.152}*100=3.8,%

----------------

If the cart did not accelerate, velocity will be constant.( acceleration is zero)

x=x_0+ut , v=u

x-t graph v-t graph